Please help me with this problem: Your supervisor instructs you to purchase 240 pens and 6 staplers for the nurse's station. Pens are purchased in sets of 6 for $2.35 per pack. Staplers are sold in sets of 2 for 12.95. How much will purchasing these products cost?

Answer by richwmiller(17219)

How many packs of pens do you need?
240/6=40 packs
40*2.35=cost of pens
How many packs of staplers do you need?
6/2=3 packs of staplers
3*12.95=cost of staplers
total cost=40*2.35+3*12.95
plus tax (and shipping and handling or gas)
But we are in Math world where no one pays for taxes, S&H nor gas